OpenStudy (freemap):

What is the equation of the circle with center at (2, 4) and radius of 6?

OpenStudy (freemap):

I didn't know how to work it out so I think thats right

OpenStudy (freemap):

wait how did you get 36. How does that equation give you 36

OpenStudy (barreraa):

like this Equation of a circle with radius 'r' and center (g, f) is given by, (x - g)² + (y - f)² = r² so, required equation of the circle => (x - 2)² + (y - 4)² = 6² => (x - 2)² + (y - 4)² = 36 OR x² + y² - 4x - 8y - 16 = 0 ---------> (Ans)

OpenStudy (freemap):

Thanks so much

OpenStudy (barreraa):

(x-2)^2 + (y-4)^2 = 36

OpenStudy (barreraa):

and try post these questions on the math section please @freemap

OpenStudy (freemap):

Oh okay thanks I didn't notice